# Mitigating Hallucinations in Large Vision-Language Models via Summary-Guided Decoding

### v1

@@ Abstract
Large Vision-Language Models (LVLMs) demonstrate impressive capabilities in generating detailed and coherent responses from visual inputs. However, they are prone to generate hallucinations due to an over-reliance on language priors. To address this issue, we investigate the language priors in LVLMs and make two key observations: (1) Even when predicting the tokens associated with image-related part-of-speech (POS), models increasingly rely on linguistic priors as the token sequences grow, thereby amplifying hallucinations. (2) Methods that directly calibrate LVLM’s output distribution to mitigate language priors can lead to a degradation in text quality or even exacerbate hallucinations. Based on these findings, we propose a novel method, Summary-Guided Decoding (SGD). This method naturally encourages the model to focus more on image information by reducing the text context through summaries, while controlling only the image-related POS tokens to maintain text quality. Through experiments, we demonstrate that SGD achieves state-of-the-art performance on object hallucination benchmarks. Furthermore, in terms of the trade-off between precision and recall, SGD achieves Pareto optimality among the existing methods. Lastly, we observe that although existing methods struggle to balance the reduction of object hallucinations with maintaining text quality, SGD demonstrates robustness in handling this challenge.
@@ 1 Introduction
Large Vision-Language Models (LVLMs) have shown remarkable advancements by integrating the reasoning capabilities of Large Language Models (LLMs) to interpret visual knowledge (Zhu et al., 2023; Dai et al., 2023; Liu et al., 2024c; Li et al., 2023a). Despite their significant utility, they suffer from a critical drawback known as object hallucination, where the model generate responses that contradict the visual input Li et al. (2023b); Liu et al. (2024b). Recent studies have shown that this occurs because LVLMs rely too heavily on learned textual patterns, which referred as language priors Zhou et al. (2024); Liu et al. (2024a); Jing et al. (2023); Lee et al. (2024a). This over-reliance on language priors tends to intensify when the model generates longer sequences or detailed descriptions Favero et al. (2024), leading to frequent hallucinations as shown in Figure 1.
@@
Next, we examine the limitations of contrastive decoding, a promising methods for mitigating hallucinations Favero et al. (2024); Wang et al. (2024); Leng et al. (2023); Kim et al. (2024); Zhu et al. (2024). Our analysis reveals two primary issues: (1) The effort to reduce language priors through contrastive decoding can disrupt the natural distribution of language-related tokens, potentially degrading overall text quality. (2) As token length increases, the model’s reliance on language priors becomes more pronounced, leading the two output distributions being contrasted to become increasingly similar. This similarity reduces the effectiveness of contrastive decoding in steering the model towards an image-aligned distribution. These findings suggest that reducing language priors may be more effectively achieved by integrating visual information naturally, with minimal intervention in the decoding process.
Building on these observations, we propose a novel method called Summary-Guided Decoding (SGD). Our approach employs a summarization technique that selectively retains essential information from previously generated sentences, encouraging LVLMs to more effectively incorporate image information. To minimize unnecessary intervention for preserving text quality, the summarization is referenced only when predicting image-related POS tokens, which require image-specific details.
Our experimental results demonstrate that SGD significantly outperforms all other decoding approaches in object hallucination benchmarks (e.g., up to +16.5% in CHAIRS and +19% in CHAIRI) across various models and architecture sizes. Additionally, SGD demonstrates Pareto optimal performance, effectively balancing the reduction of object hallucinations with the preservation of high object recall. This balance becomes more pronounced as token length increases. Finally, the results confirm that SGD not only reduces object hallucinations but also preserves the overall text quality of LVLMs.
@@
  • Based on these findings, we propose Summary-Guided Decoding (SGD). SGD modifies next-token probabilities using summarized contexts, but only for image-related POS tokens. This approach aims to reflect image information while preserving LVLM’s text quality as much as possible.
  • SGD demonstrates state-of-the-art performance in object hallucination benchmarks and achieves Pareto optimal across all methods in terms of the precision-recall trade-off. Additionally, SGD preserves text quality almost entirely.
@@ 2.2 Analysis of language priors by Part-of-Speech (POS) type
Another important observation, as shown in Figure 2 (b), is that even for image-related POS tokens (e.g., NOUN), the distributional distance decreases significantly as the token length increases. This suggests that even when image information is required during decoding, models primarily rely on textual patterns. In other words, token length (or input length) has a significant influence on how language priors are employed.
Figure 4: Illustration of our Summary-Guided Decoding.
@@ 3 Summary-Guided Decoding
Based on insights from Section 2, we identify that an increase in input length results in greater reliance on language priors, thereby exacerbating hallucinations in LVLMs. To address this, we present Summary-Guided Decoding (SGD), a novel method for controlling the length of conditioning input during decoding. In SGD, we shorten the conditioning input by summarizing the previously generated text after each sentence completion. This process preserves the critical context from earlier outputs while keeping the input concise. The summarized text, combined with the image, serves as part of the conditioned input for generating the next sentence. This approach effectively reduces the input length, allowing the model to stay more focused on the provided image.
Using summarized inputs can reduce contextual information, which may cause discrepancies with the language patterns previously learned by the model. This can result in distributional shifts that weaken the model’s language modeling capabilities, ultimately degrading the quality of the generated text. To address this, we preserve the original distribution for tokens related to language modeling, while using SGD to control only the image-related POS tokens. [^2] As shown in Figure 2, we selected PROPN, ADJ, NOUN, and NUM as image-related POS. Our method is illustrated in Figure 4.
We introduce two variations of SGD for summary model usage. The first approach leverages the instruction-following capabilities inherent in LVLMs. By providing summary instructions directly to the LVLM, this method enables the model to perform SGD without incurring additional memory costs. However, a limitation of this approach is the increased computational burden, as the LVLM generates its summaries during the process. To address these challenges, we distill the summarization capability into a smaller, more efficient model, Flan-T5-base Chung et al. (2022) (see Appendix C for details). This model significantly reduces computational overhead while maintaining the advantage of input length control. We report results for both SGD with Self-Summarization (SGD-S) and SGD with the Distilled-Flan-T5 model (SGD-D), highlighting the trade-offs between efficiency and performance.
@@ 4 Experiment
SGD-D (Ours) 42.5 11.8 77.8 42.5 12.3 72.7 43.0 10.9 77.7 44.5 11.6 69.2 43.1 11.7 74.4
SGD-S (Ours) 43.0 11.1 79.1 43.5 11.9 72.2 41.5 11.7 77.3 44.5 10.4 68.8 43.1 11.3 74.4
@@ 4.1 Experiment settings
To provide a more comprehensive assessment of hallucinations, we use the Sentence-level Hallucination Ratio (SHR) Zhao et al. (2023), a GPT-4-based evaluation metric. This metric includes hallucinations involving object existence, relationships, and attributes. We generate descriptions for 200 images from the VG dataset Krishna et al. (2016), using the same prompts as in the CHAIR metric. Specifically, SHR leverages GPT-4333We used GPT-4o (gpt-4o-2024-08-06) for hallucination judgement. to compare the model’s responses with the manually annotated descriptions from the VG dataset, evaluating each response on a sentence-by-sentence to identify potential hallucinations accurately.
@@ 4.2 Main Results
Results on CHAIR evaluation. As shown in Table 1, SGD significantly outperforms the baseline methods in the CHAIRS and CHAIRI across different model sizes and architectures. Specifically, compared to Greedy decoding, SGD-S achieves a 16.5% improvement in CHAIRS and a 19% improvement in CHAIRI on LLAVA 1.5 7B. On InstructBLIP 7B, the improvements are even more pronounced, with a 23.7% improvement in CHAIRI.
@@
We conduct the CHAIR evaluation by fixing the generated token lengths at 64, 128, 256, and 512, representing a range from short to long text generation to ensure a fair evaluation of object hallucination across different methods [^4] CHAIR is a precision-based metric, which means it can be hacked by generating shorter captions or fewer objects. (see Appendix F for full experimental results). As illustrated in Figure 5 (a) our approach maintains a Pareto optimal position in the factuality-recall trade-off relative to all other methods. Notably, this robustness in managing the trade-off becomes more pronounced as the sequence length increases. Furthermore, Figure 5 (b) shows that, even when considering object hallucination alone, our method exhibits the lowest degree of object hallucination across all variations of generated token lengths. This result is significant, as it suggests that our method can capture both factual accuracy and detailed explanations across short and long generations. This demonstrates the broad applicability of our method.
@@
Results on Sentence-level Hallucination Ratio. Table 2 shows that SGD-S achieves the lowest sentence-level hallucination rate on both the LLAVA 1.5 and InstructBLIP models. Additionally, SGD-D ranks second on both models. Based on these results, our SGD method demonstrates strong factual accuracy in holistic hallucination evaluations. OPERA performs comparably to SGD-D, but since it relies on beam search, it is less efficient than our method in terms of cost. Moreover, an examination of the Sentences Per Image (SPI) reveals that our method does not achieve favorable results simply by generating fewer sentences.
@@
5 Analysis
5.1 Analysis of SGD and Contrastive Decoding
In this section, we analyze SGD and contrastive decoding, focusing on their relationship with language priors. To explore this, we compute the JSD between each method’s output and LLM distribution at each decoding step, followed by Section 2.1. For the analysis, we generate descriptions for 200 images from the MSCOCO 2014 validation set using LLAVA 1.5 7B. Factual accuracy is evaluated using the CHAIR metric, while text quality is assessed by GPT-4o OpenAI (2024) on a 1 to 5 scale (see details in Appendix E).
@@ 5.1 Analysis of SGD and Contrastive Decoding
To investigate the effectiveness of contrastive decoding when language priors significantly influence the original distribution, we investigate VCD. In VCD, the output distribution of the original image prompt is contrasted with that of the distorted image prompt to produce outputs that more align with the original image. A noteworthy observation is that both the output distributions of the original and distorted image prompts progressively converge towards the LLM distribution, as shown in Figure 6 (b). This finding indicates that language priors are influencing both the original output distribution and the output distribution that needs to be compared. Consequently, the two distributions become increasingly similar, diminishing the effectiveness of contrastive decoding. Table 3 demonstrates the reduced effectiveness of contrastive decoding, as VCD results in more instances of object hallucination compared to greedy decoding. Although current contrastive decoding methods focus on distorting the image to create meaningful differences from the original Leng et al. (2023); Kim et al. (2024); Wan et al. (2024), the strong influence of language priors may obscure the intended effects of these distortions, undermining the effectiveness of contrastive decoding. This finding is crucial to understanding the limitations of current contrastive decoding approaches.
Unlike contrastive decoding, SGD excels at reducing object hallucinations while also maintaining a good balance in terms of text quality, as shown in Table 3. To further understand the effectiveness of SGD, we measure how much SGD and the Greedy method rely on language priors for each POS type. As seen in Figure 7, SGD demonstrates a clear reduction in language priors when predicting image-related POS tokens, while preserving the original dependency on language-related POS tokens. These results indicate that our approach effectively mitigates language priors without compromising the core language modeling properties of the LVLM.
Figure 7: Average JSD values for image-related POS and language-related POS across intervals of 32 tokens, measured in Greedy decoding and SGD.
@@
Table 4: Ablation study in terms of Summary Quality and POS Control in SGD (max new tokens is 512).
5.2 Ablation study of SGD
In this section, we conduct ablation studies to evaluate the quality of the summary and the effect of POS control in SGD. For this, we use LLAVA 1.5 7B to generate descriptions for 200 images from the MSCOCO 2014 validation dataset. We employ the CHAIR metric and the text quality metric as described in Section 5.1. Additionally, we include the n-gram fluency indicator Zhao et al. (2024), represented by $\frac{set(ngrams(s))}{len(ngram(s))}$, where $s$ denotes the description, to measure fluency.
Summary Quality. We conduct an ablation experiment to evaluate the quality of summaries used in SGD. To achieve this, we employ three distinct summarization models—Distilled-Flan-T5-base, LLAVA 1.5 7B, and GPT-4o. The results, as presented in Table 4, reveal that the effect of summarization quality is consistent across these models in terms of both CHAIR and text quality. This suggests that both SGD-D and SGD-S achieve satisfactory levels of summarization quality.
POS Control. We analyze the effect of applying image-related POS control in SGD. As shown in Table 4, applying SGD to all POS tokens, as well as selectively to image-related POS tokens, reduces object hallucination compared to the original decoding method. However, when SGD is applied to all POS tokens, text quality declines compared to the baseline, with the score dropping from 4.9 to 4.06. This decline is accompanied by a notable decrease in n-gram fluency and object recall, indicating more repetitive generation. In contrast, when SGD is applied only to image-related POS tokens, the resulting text quality remains almost unchanged, with the score only slightly decreasing from 4.9 to 4.84. These results demonstrate that applying SGD selectively to image-related POS tokens effectively preserves the model’s text quality.

### v3

@@ Abstract
Large Vision-Language Models (LVLMs) demonstrate impressive capabilities in generating detailed and coherent responses from visual inputs. However, they are prone to generate hallucinations due to an over-reliance on language priors. To address this issue, we investigate the language priors in LVLMs and make two key observations: (1) Even when predicting the tokens associated with image-related part-of-speech (POS), models increasingly rely on linguistic priors as the token sequences grow, thereby amplifying hallucinations. (2) Methods that directly calibrate LVLM’s output distribution to mitigate language priors can lead to a degradation in text quality or even exacerbate hallucinations. Based on these findings, we propose a novel method, Summary-Guided Decoding (SumGD). This method naturally encourages the model to focus more on image information by reducing the text context through summaries, while controlling only the image-related POS tokens to maintain text quality. Through experiments, we demonstrate that SumGD achieves state-of-the-art performance on object hallucination benchmarks. Furthermore, in terms of the trade-off between precision and recall, SumGD achieves Pareto optimality among the existing methods. Lastly, we observe that although existing methods struggle to balance the reduction of object hallucinations with maintaining text quality, SumGD demonstrates robustness in handling this challenge.
@@ 1 Introduction
Large Vision-Language Models (LVLMs) have shown remarkable advancements by integrating the reasoning capabilities of Large Language Models (LLMs) to interpret visual knowledge (Zhu et al., 2023; Dai et al., 2023; Liu et al., 2024c; Li et al., 2023a). Despite their significant utility, they suffer from a critical drawback known as object hallucination, where the model generate responses that contradict the visual input Li et al. (2023c); Liu et al. (2024b). Recent studies have shown that this occurs because LVLMs rely too heavily on learned textual patterns, which referred as language priors Zhou et al. (2024); Liu et al. (2024a); Jing et al. (2023); Lee et al. (2024b). This over-reliance on language priors tends to intensify when the model generates longer sequences or detailed descriptions Favero et al. (2024), leading to frequent hallucinations as shown in Figure 1.
@@
Next, we examine the limitations of contrastive decoding, a promising methods for mitigating hallucinations Favero et al. (2024); Wang et al. (2024); Leng et al. (2023); Kim et al. (2024c); Zhu et al. (2024). Our analysis reveals two primary issues: (1) The effort to reduce language priors through contrastive decoding can disrupt the natural distribution of language-related tokens, potentially degrading overall text quality. (2) As token length increases, the model’s reliance on language priors becomes more pronounced, leading the two output distributions being contrasted to become increasingly similar. This similarity reduces the effectiveness of contrastive decoding in steering the model towards an image-aligned distribution. These findings suggest that reducing language priors may be more effectively achieved by integrating visual information naturally, with minimal intervention in the decoding process.
Building on these observations, we propose a novel method called Summary-Guided Decoding (SumGD). Our approach employs a summarization technique that selectively retains essential information from previously generated sentences, encouraging LVLMs to more effectively incorporate image information. To minimize unnecessary intervention for preserving text quality, the summarization is referenced only when predicting image-related POS tokens, which require image-specific details.
Our experimental results demonstrate that SumGD significantly outperforms all other decoding approaches in object hallucination benchmarks (e.g., up to +16.5% in CHAIRS and +19% in CHAIRI) across various models and architecture sizes. Additionally, SumGD demonstrates Pareto optimal performance, effectively balancing the reduction of object hallucinations with the preservation of high object recall. This balance becomes more pronounced as token length increases. Finally, the results confirm that SumGD not only reduces object hallucinations but also preserves the overall text quality of LVLMs. [^1] The code will be available at https://github.com/andy9705/SumGD
@@
  • Based on these findings, we propose Summary-Guided Decoding (SumGD). SumGD modifies next-token probabilities using summarized contexts, but only for image-related POS tokens. This approach aims to reflect image information while preserving LVLM’s text quality as much as possible.
  • SumGD demonstrates state-of-the-art performance in object hallucination benchmarks and achieves Pareto optimal across all methods in terms of the precision-recall trade-off. Additionally, SumGD preserves text quality almost entirely.
@@ 2.2 Analysis of language priors by Part-of-Speech (POS) type
Another important observation, as shown in Figure 2 (b), is that even for image-related POS tokens (e.g., NOUN), the distributional distance decreases significantly as the token length increases.
This suggests that even when image information is required during decoding, models primarily rely on textual patterns. In other words, token length (or input length) has a significant influence on how language priors are employed.
@@ 3 Summary-Guided Decoding
Based on insights from Section 2, we identify that an increase in input length results in greater reliance on language priors, thereby exacerbating hallucinations in LVLMs. To address this, we present Summary-Guided Decoding (SumGD), a novel method for controlling the length of conditioning input during decoding. In SumGD, we shorten the conditioning input by summarizing the previously generated text after each sentence completion. This process preserves the critical context from earlier outputs while keeping the input concise. The summarized text, combined with the image, serves as part of the conditioned input for generating the next sentence. This approach effectively reduces the input length, allowing the model to stay more focused on the provided image.
Using summarized inputs can reduce contextual information, which may cause discrepancies with the language patterns previously learned by the model. This can result in distributional shifts that weaken the model’s language modeling capabilities, ultimately degrading the quality of the generated text. To address this, we preserve the original distribution for tokens related to language modeling, while using SumGD to control only the image-related POS tokens. [^3] As shown in Figure 2, we selected PROPN, ADJ, NOUN, and NUM as image-related POS. Our method is illustrated in Figure 4.
We introduce two variations of SumGD for summary model usage. The first approach leverages the instruction-following capabilities inherent in LVLMs. By providing summary instructions directly to the LVLM, this method enables the model to perform SumGD without incurring additional memory costs. However, a limitation of this approach is the increased computational burden, as the LVLM generates its summaries during the process. To address these challenges, we distill the summarization capability into a smaller, more efficient model, Flan-T5-base Chung et al. (2022) (see Appendix E for details). This model significantly reduces computational overhead while maintaining the advantage of input length control. We report results for both SumGD with Self-Summarization (SumGD-S) and SumGD with the Distilled-Flan-T5 model (SumGD-D), highlighting the trade-offs between efficiency and performance.
@@ 4 Experiment
SumGD-D (Ours) 42.5 11.8 77.8 42.5 12.3 72.7 43.0 10.9 77.7 44.5 11.6 69.2 43.1 11.7 74.4
SumGD-S (Ours) 43.0 11.1 79.1 43.5 11.9 72.2 41.5 11.7 77.3 44.5 10.4 68.8 43.1 11.3 74.4
@@ 4.1 Experiment settings
To provide a more comprehensive assessment of hallucinations, we use the Sentence-level Hallucination Ratio (SHR) Zhao et al. (2023), a GPT-4-based evaluation metric. This metric includes hallucinations involving object existence, relationships, and attributes. We generate descriptions for 200 images from the VG dataset Krishna et al. (2016), using the same prompts as in the CHAIR metric. Specifically, SHR leverages GPT-$4$We used GPT-4o (gpt-4o-2024-08-06) for hallucination judgement. to compare the model’s responses with the manually annotated descriptions from the VG dataset, evaluating each response on a sentence-by-sentence to identify potential hallucinations accurately.
@@ 4.2 Main Results
Results on CHAIR evaluation. As shown in Table 1, SumGD significantly outperforms the baseline methods in the CHAIRS and CHAIRI across different model sizes and architectures. Specifically, compared to Greedy decoding, SumGD-S achieves a 16.5% improvement in CHAIRS and a 19% improvement in CHAIRI on LLAVA 1.5 7B. On InstructBLIP 7B, the improvements are even more pronounced, with a 23.7% improvement in CHAIRI.
@@
We conduct the CHAIR evaluation by fixing the generated token lengths at 64, 128, 256, and 512, representing a range from short to long text generation to ensure a fair evaluation of object hallucination across different methods [^5] CHAIR is a precision-based metric, which means it can be hacked by generating shorter captions or fewer objects. (see Appendix H for full experimental results). As illustrated in Figure 5 (a) our approach maintains a Pareto optimal position in the factuality-recall trade-off relative to all other methods. Notably, this robustness in managing the trade-off becomes more pronounced as the sequence length increases. Furthermore, Figure 5 (b) shows that, even when considering object hallucination alone, our method exhibits the lowest degree of object hallucination across all variations of generated token lengths. This result is significant, as it suggests that our method can capture both factual accuracy and detailed explanations across short and long generations. This demonstrates the broad applicability of our method.
@@
Results on Sentence-level Hallucination Ratio. Table 2 shows that SumGD-S achieves the lowest sentence-level hallucination rate on both the LLAVA 1.5 and InstructBLIP models. Additionally, SumGD-D ranks second on both models. Based on these results, our SumGD method demonstrates strong factual accuracy in holistic hallucination evaluations. OPERA performs comparably to SumGD-D, but since it relies on beam search, it is less efficient than our method in terms of cost. Moreover, an examination of the Sentences Per Image (SPI) reveals that our method does not achieve favorable results simply by generating fewer sentences.
5 Analysis
5.1 Analysis of SumGD and Contrastive Decoding
@@
In this section, we analyze SumGD and contrastive decoding, focusing on their relationship with language priors. To explore this, we compute the JSD between each method’s output and LLM distribution at each decoding step, followed by Section 2.1. For the analysis, we generate descriptions for 200 images from the MSCOCO 2014 validation set using LLAVA 1.5 7B. Factual accuracy is evaluated using the CHAIR metric, while text quality is assessed by GPT-4o OpenAI (2024) on a 1 to 5 scale (see details in Appendix G).
@@
To investigate the effectiveness of contrastive decoding when language priors significantly influence the original distribution, we investigate VCD. In VCD, the output distribution of the original image prompt is contrasted with that of the distorted image prompt to produce outputs that more align with the original image. A noteworthy observation is that both the output distributions of the original and distorted image prompts progressively converge towards the LLM distribution, as shown in Figure 6 (b). This finding indicates that language priors are influencing both the original output distribution and the output distribution that needs to be compared. Consequently, the two distributions become increasingly similar, diminishing the effectiveness of contrastive decoding. Table 3 demonstrates the reduced effectiveness of contrastive decoding, as VCD results in more instances of object hallucination compared to greedy decoding. Although current contrastive decoding methods focus on distorting the image to create meaningful differences from the original Leng et al. (2023); Kim et al. (2024c); Wan et al. (2024), the strong influence of language priors may obscure the intended effects of these distortions, undermining the effectiveness of contrastive decoding. This finding is crucial to understanding the limitations of current contrastive decoding approaches.
Unlike contrastive decoding, SumGD excels at reducing object hallucinations while also maintaining a good balance in terms of text quality, as shown in Table 3. To further understand the effectiveness of SumGD, we measure how much SumGD and the Greedy method rely on language priors for each POS type. As seen in Figure 7, SumGD demonstrates a clear reduction in language priors when predicting image-related POS tokens, while preserving the original dependency on language-related POS tokens. These results indicate that our approach effectively mitigates language priors without compromising the core language modeling properties of the LVLM.
Figure 7: Average JSD values for image-related POS and language-related POS across intervals of 32 tokens, measured in Greedy decoding and SumGD.
@@
Table 4: Ablation study in terms of Summary Quality and POS Control in SumGD (max new tokens is 512).
5.2 Ablation study of SumGD
In this section, we conduct ablation studies to evaluate the quality of the summary and the effect of POS control in SumGD. For this, we use LLAVA 1.5 7B to generate descriptions for 200 images from the MSCOCO 2014 validation dataset. We employ the CHAIR metric and the text quality metric as described in Section 5.1. Additionally, we include the n-gram fluency indicator Zhao et al. (2024), represented by $\frac{set(ngrams(s))}{len(ngram(s))}$, where $s$ denotes the description, to measure fluency.
Summary Quality. We conduct an ablation experiment to evaluate the quality of summaries used in SumGD. To achieve this, we employ three distinct summarization models—Distilled-Flan-T5-base, LLAVA 1.5 7B, and GPT-4o. The results, as presented in Table 4, reveal that the effect of summarization quality is consistent across these models in terms of both CHAIR and text quality. This suggests that both SumGD-D and SumGD-S achieve satisfactory levels of summarization quality.
POS Control. We analyze the effect of applying image-related POS control in SumGD. As shown in Table 4, applying SumGD to all POS tokens, as well as selectively to image-related POS tokens, reduces object hallucination compared to the original decoding method. However, when SumGD is applied to all POS tokens, text quality declines compared to the baseline, with the score dropping from 4.9 to 4.06. This decline is accompanied by a notable decrease in n-gram fluency and object recall, indicating more repetitive generation. In contrast, when SumGD is applied only to image-related POS tokens, the resulting text quality remains almost unchanged, with the score only slightly decreasing from 4.9 to 4.84. These results demonstrate that applying SumGD selectively to image-related POS tokens effectively preserves the model’s text quality.